OpenStudy (anonymous):

D= vit +1/2at^2 Solve for vi the solve for a

OpenStudy (anonymous):

\[D=v_it+\frac{1}{2}at^2\] solving for vi start by subtracting 1/2at^2 from both sides \[D-\frac{1}{2}at^2=v_\it+\frac{1}{2}at^2-\frac{1}{2}at^2\] \[D-\frac{1}{2}at^2=v_it\] then divide t from both sides

OpenStudy (anonymous):

Ok then what

OpenStudy (anonymous):

then you've solved for v_i do the same kind of process for solving for a when being asked to solve for any variables, it means you need to isolate that variable so its by itself on one side. just remember if you ever do something to one side, you always have to do it to the other

OpenStudy (anonymous):

Do I divide t on both sides

OpenStudy (anonymous):

for which problem? v_i?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

yes you need to divide by t to both sides, that way v_i will be by itself

OpenStudy (anonymous):

So will it be d-.5at2/t=vi

OpenStudy (anonymous):

@Com

OpenStudy (anonymous):

@completeidiot

OpenStudy (anonymous):

Is my Nswer above sorrect

OpenStudy (anonymous):

Correct And answer

OpenStudy (anonymous):

it will look like \[\frac{D-\frac{1}{2} at^2}{t}=v_i\] t is divided over the whole thing

OpenStudy (anonymous):

What about the second one could you guide me through that 2

OpenStudy (anonymous):

ok, we need to solve for a so that means you should subtract v_it from both sides

OpenStudy (anonymous):

do you understand why we're subtracting v_i*t ?

OpenStudy (anonymous):

Yes to get a alone

OpenStudy (anonymous):

ok then what will you do to "a"?

OpenStudy (anonymous):

Nothing yet , I'm not sure

OpenStudy (anonymous):

ok you have 1/2* a* t^2 we need to get "a" alone what could you multiply by or divide by to get a alone

OpenStudy (anonymous):

What. Does the asters mean